OpenStudy (anonymous):

if you write 14 of the power of 6 in expandes forma ,hiw many factores will there be

OpenStudy (cwrw238):

2^3 = 2*2*2

OpenStudy (cwrw238):

14^6 = ?

OpenStudy (anonymous):

If you write 14^6 in expanded form , how many factors will there be